# Accelerating ViT Inference on FPGA through Static and Dynamic Pruning

### 21 Mar 2024 v1

@@ VI Implementation Details
Evaluated Model: We evaluate our approach on the widely used DeiT-Small [44] model, which has 12 layers with each layer having 6 heads. The hidden dimension is $D=384$ and the (base) model has $22$M parameters.
@@
Implementation details of weight pruning, token pruning, and simultaneous training: The DeiT-Small model is simultaneously pruned as per algorithm 1. We train several variants of the model by varying the model pruning top-$k$ rate $r_{b}$, token pruning keep rate $r_{t}$ and block size $b$. Specifically, $r_{b}$ is varied over $\{0.5,0.7\}$, $r_{t}$ over $\{0.5,0.7,0.9\}$ and $b$ over $\{16,32\}$. A cubic sparsity scheduler, as in [17], is used to schedule $r_{b}$ from a full density of $1$, to its final density ($0.5$ or $0.7$) with a warm-up and a cool-down phase. The token dropping module, TDM, is inserted in the $3^{\text{rd}}$, $7^{\text{th}}$ and $10^{\text{th}}$ encoder layers. All model variants are trained for a total of $30$ epochs using the AdamW optimizer [45] with a learning rate of $2\times 10^{-5}$, weight decay of $0.01$ and a batch size of $128$ across $4$ GPUs. Note that to train all pruned variants as well as the baseline, we use the pre-trained DeiT-Small model available at [46] with the classification MLP head parameters re-initialized. A ViT base model is used as the teacher for knowledge distillation.
@@
Hardware implementation details: We implement our FPGA design on a state-of-the-art FPGA platform, Xilinx Alveo U250, which consists of four Super Logic Regions (SLRs). We implement the proposed hardware design using Xilinx High-level Synthesis (HLS). For the MPCA module, we empirically determine the hardware hyperparameters to be $p_{h}=4$, $p_{t}=12$, $p_{c}=2$, $p_{pe}=8$ according to the hardware resources of the target FPGA board: (1) We set $p_{h}=4$ because Alveo u250 has four SLRs, with each SLR placed one CHM. (2) We set $p_{c}=2$ because in a CHM, the PEs of the same row load the same rows of tokens but different data blocks. The BRAM/URAM on FPGA has two independent memory ports, which can support concurrent memory access of $2$ columns of PEs ($p_{c}=2$). (3) We set $p_{pe}=8$ because the data block size $b$ is set as $16$ or $32$ for block-wise weight pruning. Using $p_{pe}=8$ can support these two block sizes without data padding as well as keeping a reasonable value for $p_{t}$. (4) For setting $p_{t}$, the PEs of the same column within a CHM shares the same weight blocks. The weight blocks are broadcast into each PE of the same column, which supports any value of $p_{t}$. We set $p_{t}$ according to the available resources of the target FPGA board after determining $p_{h}$, $p_{c}$, and $p_{pe}$. We use the int16 data format. We utilize the four DDR4 channels of U250, which have 77 GB/s of external memory bandwidth in total. We perform synthesis and place-route for the design using Xilinx Vitis v2022.2. We report the frequency and FPGA resource utilization after place-route. The achieved frequency is $300$ Mhz, and the resource utilization is shown in Table IV.
@@
LUTs DSPs URAMs BRAMs Used 798K 7088 1728 960 Available 1726K 12288 2688 1280 Utilization (%) 46% 57% 64% 75%
@@ VII-A Baselines, Metrics, Datasets
Performance Metrics: We utilize the following performance metrics: (1) Accuracy: Following prior works, we evaluate the accuracy of our pruned model on ImageNet. (2) Inference latency: Following prior works [37, 47, 35], inference latency, which is the time elapsed from the time accelerator begins the execution to the time it produces the final inference result. (3) Throughput: Throughput denotes the number of images that can be processed for a given time frame. (4) Computation complexity (FLOPS): We measure the computational complexity which is the number of floating-point operations (FLOPs). (5) Model size: The amount of memory space (MB) to store the model.
@@ VII-B Evaluation for the Pruning Algorithm
Results in Table VI indicate that for extreme pruning settings ($r_{b}$, $r_{t}$ both $0.5$), the accuracy drop ($\approx 12\%$) compared against the baseline is not insignificant. A major reason for this drop is due to the the fact that the training epochs for our experiments were restricted to $30$ despite the reduction in model and input density. With smaller top-$k$ rate $r_{b}$ and token keep rate $r_{t}$, the model requires larger epochs to converge. Compared to the baseline DeiT-Small model, the proposed simultaneous pruning algorithm achieves a compression ratio of upto 1.24×1.24$\times$1.24 × to 1.60×1.60$\times$1.60 × and a reduction in computational cost of upto 1.43×1.43$\times$1.43 × to 3.42×3.42$\times$3.42 × with an accuracy drop of as little as $\approx 3\%$. Whilst prior works focus on either reducing the model size [43] or on reducing the computational complexity [37, 35], our proposed simultaneous pruning algorithm targets both.
@@ VII-C1 Cross platform comparison
We compare the latency and throughput for executing the pruned model with baseline CPU and GPU (Figure 9 and 10). The latency of our accelerator is measured when batch size is $1$, and the throughput is calculated by $\frac{1}{\text{latency}}$. For comparing the latency, we set the batch size as $1$ for CPU and GPU because a larger batch size will increase the latency for CPU and GPU. For comparing throughput, we set the batch size as $8$ for CPU and GPU, which can fully exploit their thread-level parallelism. On average, our FPGA accelerator achieves latency reduction of 12.8×12.8$\times$12.8 × and 3.2×3.2$\times$3.2 × compared with CPU and GPU, respectively. The lower latency of our FPGA accelerator is because (1) our MPCA module with load balance strategy fully exploits the computation parallelism within the pruned model. The pruning ratios are higher ($r_{b}$ and $r_{t}$ are smaller), and our FPGA accelerator achieves higher speedup. In contrast, the CPU and GPU cannot efficiently deal with the computational irregularity caused by weight pruning and dynamic token dropping. (2) CPU and GPU have complex cache hierarchies, leading to higher memory access latency for executing ViT inference, leading to increased latency. On average, our FPGA accelerator achieves 3.6×3.6$\times$3.6 × and 0.45×0.45$\times$0.45 × throughput speedup compared with CPU and GPU, respectively. Our FPGA accelerator achieves lower throughput (0.45×0.45$\times$0.45 ×) than GPU because GPU has much higher peak performance (50×50$\times$50 ×) and eternal memory bandwidth. When the pruning ratios become high (e.g., $r_{b}=0.5$ and $r_{t}=0.5$), our throughput gets closer to GPU, which indicates our FPGA accelerator has higher efficiency for executing the ViT model with larger pruning ratio.
@@ VIII Conclusion and Future Work
In this paper, we proposed an algorithm-hardware codesign that simultaneously utilizes the static weight pruning and dynamic token pruning approaches, bridging the gap of prior works that utilize only one pruning algorithm, further reducing the computation complexity of ViT. The proposed hardware accelerator can efficiently execute the pruned model through novel hardware architecture design. In the future, we plan to develop a design automation framework that automatically generates optimized implementation for the pruned ViT model given a target FPGA platform.
@@ Acknowledgement
This work is supported by the DEVCOM Army Research Lab (ARL) under grants W911NF2220159, and the National Science Foundation (NSF) under grants SaTC-2104264. Equipment and support by AMD AECG are greatly appreciated.
Distribution Statement A: Approved for public release. Distribution is unlimited.

### 12 Apr 2024 v2

@@ VI Implementation Details
Evaluated Model: We evaluate our approach on the widely used DeiT-Small [45] model, which has 12 layers, with each layer having six heads. The hidden dimension is $D=384$, and the (base) model has $22$M parameters.
@@
Implementation details of weight pruning, token pruning, and simultaneous training: The DeiT-Small model is simultaneously pruned as per algorithm 1. We train several variants of the model by varying the model pruning top-$k$ rate $r_{b}$, token pruning keep rate $r_{t}$, and block size $b$. Specifically, $r_{b}$ is varied over $\{0.5,0.7\}$, $r_{t}$ over $\{0.5,0.7,0.9\}$ and $b$ over $\{16,32\}$. A cubic sparsity scheduler, as in [17], is used to schedule $r_{b}$ from a full density of $1$ to its final density ($0.5$ or $0.7$) with a warm-up and a cool-down phase. The token-dropping module, TDM, is inserted in the $3^{\text{rd}}$, $7^{\text{th}}$ and $10^{\text{th}}$ encoder layers. All model variants are trained for a total of $30$ epochs using the AdamW optimizer [46] with a learning rate of $2\times 10^{-5}$, weight decay of $0.01$ and a batch size of $128$ across $4$ GPUs. Note that to train all pruned variants as well as the baseline, we use the pre-trained DeiT-Small model available at [47] with the classification MLP head parameters re-initialized. A ViT base model is used as the teacher for knowledge distillation.
@@
Hardware implementation details: We implement our FPGA design on a state-of-the-art FPGA platform, Xilinx Alveo U250, which consists of four Super Logic Regions (SLRs). We implement the proposed hardware design using Xilinx High-level Synthesis (HLS). For the MPCA module, we empirically determine the hardware hyperparameters to be $p_{h}=4$, $p_{t}=12$, $p_{c}=2$, $p_{pe}=8$ according to the hardware resources of the target FPGA board: (1) We set $p_{h}=4$ because Alveo u250 has four SLRs, with each SLR placed in one CHM. (2) We set $p_{c}=2$ because in a CHM, the PEs of the same row load the same rows of tokens but different data blocks. The BRAM/URAM on FPGA has two independent memory ports, which can support concurrent memory access of $2$ columns of PEs ($p_{c}=2$). (3) We set $p_{pe}=8$ because the data block size $b$ is set as $16$ or $32$ for block-wise weight pruning. Using $p_{pe}=8$ can support these two block sizes without data padding as well as keeping a reasonable value for $p_{t}$. (4) For setting $p_{t}$, the PEs of the same column within a CHM shares the same weight blocks. The weight blocks are broadcast into each PE of the same column, which supports any value of $p_{t}$. We set $p_{t}$ according to the available resources of the target FPGA board after determining $p_{h}$, $p_{c}$, and $p_{pe}$. We use the int16 data format. We utilize the four DDR4 channels of U250, which have 77 GB/s of external memory bandwidth in total. We perform synthesis and place-route for the design using Xilinx Vitis v2022.2. We report the frequency and FPGA resource utilization after place-route. The achieved frequency is $300$ MHz, and the resource utilization is shown in Table IV.
@@
LUTs DSPs URAMs BRAMs HeatViT [37] 137.6K161.4K 1955-2066 N/A 338-528 Auto-Vit-Acc [48] 120K-193K 13-2066 N/A N/A Our Work 798K 7088 1728 960
@@ VII-A Baselines, Metrics, Datasets
Performance Metrics: We utilize the following performance metrics: (1) Accuracy: Following prior works, we evaluate the accuracy of our pruned model on ImageNet. (2) Inference latency: Following prior works [37, 49, 35], we measured inference latency via hardware emulation using AMD-Xilinx Vitis, which accurately simulates the behavior of FPGA DDR. The measured latency is end-to-end from the time when the input is given at DDR to the time when the inference result is written back to DDR. (3) Throughput: Throughput denotes the number of images that can be processed for a given time frame. (4) Computation complexity (FLOPS): We measure the computational complexity, which is the number of floating-point operations (FLOPs). (5) Model size: The amount of memory space (MB) to store the model.
@@ VII-B Evaluation for the Pruning Algorithm
Results in Table VI indicate that for extreme pruning settings ($r_{b}$, $r_{t}$ both $0.5$), the accuracy drop ($\approx 12\%$) compared against the baseline is not insignificant. A major reason for this drop is the fact that the training epochs for our experiments were restricted to $30$ despite the reduction in model and input density. With a lower top-$k$ rate $r_{b}$ and token keep rate $r_{t}$, the model requires larger epochs to converge. Compared to the baseline DeiT-Small model, the proposed simultaneous pruning algorithm achieves a compression ratio of up to 1.24×1.24$\times$1.24 × to 1.60×1.60$\times$1.60 × and a reduction in the computational cost of up to 1.43×1.43$\times$1.43 × to 3.42×3.42$\times$3.42 × with an accuracy drop of as little as $\approx 3\%$. Whilst prior works focus on either reducing the model size [44] or on reducing the computational complexity [37, 35], our proposed simultaneous pruning algorithm targets both.
@@ VII-C1 Cross platform comparison
We compare the latency and throughput for executing the pruned model with baseline CPU and GPU (Figure 9 and 10). The latency of our accelerator is measured when batch size is $1$, and the throughput is calculated by $\frac{1}{\text{latency}}$. For comparing the latency, we set the batch size as $1$ for CPU and GPU because a larger batch size will increase the latency for CPU and GPU. For throughput comparison, we set the batch size as $8$ for CPU and GPU, which can fully exploit their thread-level parallelism. On average, our FPGA accelerator achieves a latency reduction of 12.8×12.8$\times$12.8 × and 3.2×3.2$\times$3.2 ×, compared with CPU and GPU, respectively. The lower latency of our FPGA accelerator is due to the followings: (1) our MPCA module with load balance strategy fully exploits the computation parallelism within the pruned model. The FPGA accelerator achieves a higher speedup with higher pruning ratios (smaller $r_{b}$ and $r_{t}$). In contrast, the CPU and GPU cannot efficiently handle the computational irregularity caused by weight pruning and dynamic token dropping. (2) CPU and GPU have complex cache hierarchies, leading to higher memory access latency for executing ViT inference, leading to increased latency. On average, our FPGA accelerator achieves 3.6×3.6$\times$3.6 × and 0.45×0.45$\times$0.45 × throughput speedup compared with CPU and GPU, respectively. Our FPGA accelerator achieves a lower throughput (0.45×0.45$\times$0.45 ×) than GPU, because GPU has much higher peak performance (50×50$\times$50 ×) and eternal memory bandwidth. When the pruning ratios become high (e.g., $r_{b}=0.5$ and $r_{t}=0.5$), our throughput gets closer to GPU, which indicates that our FPGA accelerator has higher efficiency for executing the ViT model with larger pruning ratio.
@@ VIII Conclusion and Future Work
In this paper, we proposed an algorithm-hardware codesign that simultaneously utilizes the static weight pruning and dynamic token pruning approaches. It bridges the gap of prior works that utilize only one pruning algorithm, further reducing the computation complexity of ViT. The proposed hardware accelerator can efficiently execute the pruned model through novel hardware architecture design. In the future, we plan to develop a design automation framework that automatically generates optimized implementation for the pruned ViT model given a target FPGA platform.
@@ Acknowledgement
This work is supported by the DEVCOM Army Research Lab (ARL) under grants W911NF2220159, and the National Science Foundation (NSF) under grants CCF-1919289 and SaTC-2104264. Equipment and support by AMD AECG are greatly appreciated. Distribution Statement A: Approved for public release. Distribution is unlimited.